After a two weeks of an ODI hiatus, we are back at the telescope. We allowed ODI to warm up for some maintenance work, and today we spend most time on pumping the dewar and going through the cool down procedure.The heater to maint a stable temperature of 170ºK just kicked in an hour ago.
Andrey is spending the time to apply some software updates to further improve the capturing of telemetry data and to enable automated data transfer to the ODI Pipeline, Portal, and Archive (PPA). We will test the new updates before the end of the week to ensure we really made things ..... "better". However, we are somewhat limited in testing since we can operate only during the daytime. By 2 to 3pm we have to yield to the regular observers at WIYN.
Next week we will have the first regular scientific commissioning run for ODI, which is a change in our operating mode: Now observers from the WIYN community, but outside the ODI team, take over and help with the performance characterization of ODI. This will also be an exciting time for all of us.
I finally managed to get the full ODI installation time lapse video online; Pete Marenfeld edited this final version, and it looks really good. It is hard to imagine that more than two month have passed since that movie was recorded.
ODI-installation15FPS from Daniel Harbeck on Vimeo.